A 48.8-mL sample of a 7.8 M KNO3 solution is diluted to 1.20 L .

1.What volume of the diluted solution contains 10.0 g of KNO3? (Hint: Figure out the concentration of the diluted solution first.)Express your answer using two significant figures.

2. To what volume should you dilute 130 mL of an 8.00 M CuCl2 solution so that 50.0 mL of the diluted solution contains 5.7 g CuCl2?

Express your answer using two significant figures.
Answer #1

dilute solution                 concentrated solution

M1 =                                     M2 = 7.8M

V1 = 1.2L = 1200ml               V2 = 48.8ml

               M1V1   = M2V2

               M1     = M2V2/V1

                          = 7.8*48.8/1200

                            = 0.3172M

no of moles of KNO3   = W/G.M.Wt

                                     = 10/101   = 0.099 moles

molarity   = no of moles/volume in L

0.3172         = 0.099/volume in L

volume in L   = 0.099/0.3172   = 0.31L

volume of KNO3   = 310ml

part-B

no of moles of CuCl2   = W/G.m.Wt

                                     = 5.7/134.452   = 0.0424moles

molarity   = no of moles/volume in L

              = 0.0424/0.05   = 0.848M

diluted solution                  concentrated solution

M1 = 0.848M                      M2 = 8M

V1   =                                    V2 = 130ml

                 M1V1   = M2V2

                   V1      =   M2V2/M1

                              = 8*130/0.848

                             = 1226.4ml

                              = 1.2*10^3ml >>>>answer